OpenStudy (anonymous):

s=(1/3)gt^2 solve for t

OpenStudy (australopithecus):

s/((1/3)g) = t^(2) (s/(1/3)g))^(1/2) = t^(2)^(1/2) (s/(1/3)g))^(1/2) = t^(2/2) (s/(1/3)g))^(1/2) = t^(1) = t

OpenStudy (kropot72):

\[s=(1/3)g \times t ^{2}\] Dividing both sides by (1/3)g gives: \[t ^{2}=3s \div g\] \[t=\sqrt{(3s)\div g}\]

OpenStudy (australopithecus):

s/((1/3)g) = t^(2) (s/(1/3)g))^(1)(1/2) = t^(2)(1/2) (s/(1/3)g))^(1/2) = t^(2/2) (s/(1/3)g))^(1/2) = t^(1) = t

OpenStudy (anonymous):

thank you all

OpenStudy (australopithecus):

wouldn't it equal |t|

OpenStudy (australopithecus):

as t cannot be a negative number